OpenStudy (anonymous):

finding the slope of an equation of a line. y=-10x-5

OpenStudy (anonymous):

y=mx+b m=slope y=-10x-5 .......^ slope=10

OpenStudy (yanasidlinskiy):

y=mx+b y= -10+5

OpenStudy (anonymous):

yana x+b ..........^ doesn't have to be a plus, can be y=-10x-5

OpenStudy (yanasidlinskiy):

Yea. I know that. That is just the formula to find the slope. -10x is the slope and the -5 is b

OpenStudy (yanasidlinskiy):

Kaylaa. Do you understand how to find the slope?

OpenStudy (anonymous):

no the -10 is the slope, the x is extra y=(m)x+b m alone is the slope without the x

OpenStudy (yanasidlinskiy):

Ok. Well that's what I meant.